OpenStudy (anonymous):

factor any method. 33x^2+7x-10

OpenStudy (anonymous):

(3x+2)(11x−5)

OpenStudy (anonymous):

\[33x^2+7x-10\] 33*-10=-330 22-15=7 22*-15=-330 write 7x=22x-15x and then take common to make factors.

OpenStudy (anonymous):

youvan make factors using quadratic formula. consider \[33x^2+7x-10=0,x=\frac{ -7\pm \sqrt{7^2-4*33*-10} }{ 2*33 }\] \[x=\frac{ -7\pm \sqrt{49+1320} }{66}=\frac{ -7\pm \sqrt{1369} }{ 66 }=\frac{ -7\pm37 }{ 66}\] \[x=\frac{ 30 }{66 }=\frac{ 5 }{11 },hence~ x-\frac{ 5 }{ 11 } ~is~a~factor\] or 11x-5 is a factor. or \[x=\frac{ -44 }{ 66 }=\frac{ -2 }{3 },or~x+\frac{ 2 }{3 }~is~a~factor.\] or 3x+2 is a factor. \[33x^2+7x-10=\left( 11x-5 \right)\left( 3x+2 \right)\]